# CzechLynx: A Dataset for Individual Identification and Pose Estimation of the Eurasian Lynx

### v1

@@ Abstract
We introduce CzechLynx, the first large-scale, open-access dataset for individual identification, 2D pose estimation, and instance segmentation of the Eurasian lynx (Lynx lynx). CzechLynx includes more than 30k camera trap images annotated with segmentation masks, identity labels, and 20-point skeletons and covers 219 unique individuals across 15 years of systematic monitoring in two geographically distinct regions: Southwest Bohemia and the Western Carpathians. To increase the data variability, we create a complementary synthetic set with more than 100k photorealistic images generated via a Unity-based pipeline and diffusion-driven text-to-texture modeling, covering diverse environments, poses, and coat-pattern variations. To allow testing generalization across spatial and temporal domains, we define three tailored evaluation protocols/splits: (i) geo-aware, (ii) time-aware open-set, and (iii) time-aware closed-set. This dataset is targeted to be instrumental in benchmarking state-of-the-art models and the development of novel methods for not just individual animal re-identification.
@@ 1 Background & Summary
Understanding species movement, population structure, and habitat connectivity is critical for effective conservation, especially for wide-ranging, protected carnivores. In Europe, the Eurasian lynx (Lynx lynx) exemplifies these challenges, being ecologically important, legally protected, and indicative of broader biodiversity concerns. This medium-sized, solitary, and territorial carnivore lives at low population densities, with substantial home ranges estimated at 282.9$\pm$42.4 km2 for males and 147.6$\pm$50.3 km2 for females in Central Europe [1]. Although primarily inhabiting forested areas and rugged terrains, the lynx demonstrates notable adaptability by using refuge areas to mitigate human pressures [2] and can disperse across Central Europe’s cultural landscapes, covering distances greater than 98 km from source populations [3]. Extensive and long-term monitoring efforts through camera-trapping have generated rich datasets, valuable not only for ecological research, such as estimating population densities and demographic parameters [4, 5, 6, 7], but also for the development and evaluation of novel machine learning methods. These datasets support progress in computer vision tasks, including fine-grained recognition, individual re-identification, and open-set recognition, enhancing the accuracy and efficiency of wildlife monitoring. This creates a mutually beneficial feedback loop: ecological data fuels machine learning research, and machine learning makes processing raw conservation data easier. Consolidating and standardizing these datasets for machine learning unlocks their full potential for both domains. Several studies have already been published based on the available data.
Straka et al. [8] used part of CzechLynx dataset to develop methods for 2D pose estimation tailored to endangered species. Their study focused on addressing the challenges of limited data availability by combining real images from our Central European lynx monitoring project with synthetically generated data. To ensure a robust cross-regional evaluation, they constructed training and testing datasets from two geographically distinct regions and two lynx populations: Southwest Bohemia (a Czech, Bavarian, and Austrian population) and the Western Carpathians (a Carpathian population). Using our dataset, they fine-tuned and benchmarked various state-of-the-art models, including HRNet-w32[9], and demonstrated that incorporating real lynx data significantly improved pose estimation accuracy. Their work highlights the importance and effectiveness of high-quality, species-specific datasets like ours in advancing computer vision techniques for wildlife conservation.
@@
Picek et al. [10] used the CzechLynx dataset to evaluate a method for individual animal identification, which models foreground and background information separately. They focused on the challenges of re-identifying lynx individuals over time and across different locations, using images from the long-term camera-trap monitoring project in Central Europe. The dataset enabled them to demonstrate how incorporating spatial and temporal priors, along with their Per-Instance Temperature Scaling (PITS), improves identification accuracy, particularly when animals appear in new areas or exhibit changes in appearance over time.
@@
Dula et al. [6] and Palmero et al. [7] used some parts of the CzechLynx dataset to estimate population densities by spatial capture–recapture (SCR) models and other demographic parameters of the Eurasian lynx in the West Carpathians in the Czech-Slovakia borderland and in the Bohemian Forest Ecosystem: Bavarian Forest National Park in Germany and the Šumava National Park in the Czech Republic. By identifying lynx individuals, Dula et al. [6] estimated apparent survival, transition rates, and population turnover. Fluctuating densities and high turnover rates indicated human-caused mortality, which could limit population growth and the dispersal of lynx to other adjacent areas, thus undermining the favourable conservation status of the Carpathian population.
@@ 1.1 Related Work
Existing animal-focused datasets cover a variety of species and provide foundational resources for developing and evaluating methods for individual animal identification. However, there is a geographic bias, with most publicly accessible datasets originating from North America and Africa, while European datasets remain less commonly shared due to data privacy and conservation concerns. Recent initiatives have aimed to create datasets specifically for the identification of individuals [11]. Currently available datasets include a wide array of mammals (e.g., zebras, belugas, whale sharks, and elephants), carnivores (e.g., leopards, hyenas, and tigers), and reptiles (e.g., turtles). However, datasets featuring large carnivores listed on the EU red lists are notably underrepresented. A summary of publicly available datasets and their relevant statistics is provided in Table 2, and non-public datasets are listed in Table 1. Brief descriptions of selected publicly accessible datasets are included below.
@@
Table 1: Non-public animal re-id datasets.
The ATRW dataset [20] comprises 8,076 videos collected from ten Zoos in China and covers 92 Amur tigers (Panthera tigris). Each image is annotated with bounding boxes, identity labels, and pose keypoints, allowing development of re-identification and pose estimation methods for controlled environments.
The Hyena ID 2022 and Leopard ID 2022 datasets [21], developed by the Botswana Predator Conservation Trust, contain 3,129 images of 256 individual spotted hyenas (Crocuta crocuta) and 6,795 images of 430 African leopards (Panthera pardus), respectively. Images were taken in natural settings and annotated with bounding boxes and viewpoints, making these datasets valuable for testing in-the-wild generalization performance for individual animal re-id.
The LionData dataset [22] was compiled as part of the Mara Masai project in Kenya and includes 750 images of 98 lions (Panthera leo). Each image was manually annotated to support training and evaluation of re-id models under natural variability, such as lighting changes and different camera angles.
@@
Name Species Records Individuals Span BBoxes Masks Pose In-the-wild
@@
ATRW [20] Pantheria tigris 8,076   92 1 day $\bullet$ $\bullet$ $\circ$ $\circ$
CzechLynx (our) Lynx lynx 37,440 219 15 years $\bullet$ $\bullet$ $\bullet$ $\bullet$
Table 2: All publicly available datasets with large carnivores. datasets based on video recordings. excluding synthetic data.
@@ 2 Methods
This section outlines the procedures used to create and prepare the CzechLynx dataset, including detailed descriptions of data acquisition, annotation, and preprocessing steps. The dataset combines data collected from long-term camera trapping surveys across two regions with synthetic samples generated to enhance variability. Where applicable, previously published methods are cited, and only novel technical processes are described in detail.
@@ 2.1 Dataset Collection
The CzechLynx dataset comprises data collected from two independent conservation projects operating within distinct regions of the Eurasian lynx’s (Lynx lynx) Central European range. Each project used its own field methodology and data management pipeline. To maintain transparency and reproducibility, we describe the sources separately below.
@@
Carnivore Conservation Programme, Friends of the Earth Czech Republic (FoE CZ), and Department of Forest Ecology, Mendel University in Brno: This source includes data collected through long-term monitoring in the Western Carpathians (since 2009) and Southwest Bohemia (since 2015). The datasets are labeled "FoE CZ - The Western Carpathians" and "FoE CZ - Southwest Bohemia" in the following text. Camera traps were deployed on forest roads, trails, mountain ridges, as well as marking and resting sites, identified via snow-tracking, signs of occurrence, or telemetry research [24, 1]. From 2015 onward, the Western Carpathians area followed a systematic camera-trapping and spatial capture-recapture model to estimate lynx densities and evaluate density fluctuations, apparent survival, transition rate, and individual turnover during five consecutive seasons [6]. Earlier data and those collected by opportunistic camera-trapping allowed for estimation of minimum population size, social structure, and evidence of reproduction. Cameras used included both white-flash and infrared models (video-enabled). Most individuals were manually identified based on distinctive coat patterns visible on their flanks, forelimbs, and hindlimbs [6].
@@
Šumava National Park: This source includes data collected from the areas of Šumava National Park and the Protected Landscape Area in southwestern Bohemia. The data were gathered through two long-term monitoring projects conducted in cooperation with the Šumava National Park Administration. The first project, started in 2009, covers roughly the northwestern two-thirds of Šumava National Park. Its main goal is to obtain yearly density estimates of independent lynx individuals living in the core area of the Bohemian-Bavarian-Austrian lynx population. Within the scope of the first project, data have been collected using a 2.7$\times$2.7 km grid. White-flash camera traps were installed seasonally in every second grid cell at suitable locations along forest paths, roads, and trails. Additional details about this methodology can be found in studies by Weingarth et al. [25] and Palmero et al. [26]. The second project (initiated in 2017) aims to provide as complete an overview as possible of the entire Bohemian-Bavarian-Austrian lynx population throughout its full range. Data for this project were collected year-round at various sites, including forest paths, roads, trails, and lynx scent-marking locations, using both white-flash and infrared video-enabled camera traps. These cameras were set up continuously based on a 10$\times$10 km grid (ETRS LAEA 5210). In each grid cell predominantly within the Šumava National Park and Protected Landscape Area, between 4 to 8 camera sites were continuously monitored throughout the year. Further information about this project is available in a separate report [27].
@@ 2.2 Raw Data Pre-processing
The pre-processing of raw camera trap data began with the manual removal of empty images, which frequently resulted from false triggers caused by wind, moving vegetation, or fluctuating light conditions. Once these non-informative images were discarded, the remaining photographs were screened by a team of experts and trained volunteers, who classified the content by species, including instances involving humans or vehicles. Only images positively identified as containing Eurasian lynxes were retained for further analysis. Selected images underwent a second round of expert review by specialists with extensive experience in long-term lynx monitoring. Using the species’ distinctive coat patterns, especially on the hind limbs, forelimbs, and flanks, experts manually identified individual lynxes. In each study area (Southwest Bohemia and the Western Carpathians), at least three well-trained local observers conducted this detailed identification process, supported by a cross-verification aligned with established camera-trapping standards [6, 28]. The final set of verified images, each linked to a specific, individually recognized lynx, provided the essential foundation for all subsequent annotation and analysis steps.
@@ 2.3 Video Pre-processing
The CzechLynx dataset includes data collected from various camera trap models, some of which capture still images while others record short video sequences during detection events. Still cameras typically produce image bursts, usually three frames per trigger, while video-capable traps generate longer recordings. Since manually annotating every frame in a video is not feasible, we implemented a method to automatically extract a small number of informative frames from each clip.
The selection process is based on automated detection using MegaDetector [29], a YOLOv5-based object detection model. Each frame of the video is analyzed, and those with a detection confidence below 0.7 are discarded. From the remaining frames, the first and last with valid detections are identified, and the interval between them is divided into three equal parts. One frame is then selected from each segment – specifically, the frame with the largest detected bounding box – ensuring that the chosen images capture the animal prominently and are distributed across the video timeline. Selected frames serve as representative samples for downstream tasks such as individual identification and pose estimation. The following criteria guided the frame selection process: (i) the animal’s full body should be visible; (ii) the animal should be captured from a side view, where identifying coat patterns are most apparent; (iii) the animal should be close enough to the camera to reveal fine details without any part being cropped; and (iv) the selected frames should be temporally spaced to capture variation in pose and movement. Figure 1 illustrates the selected frames and the relative bounding box areas across the video sequence.
Figure 1: Example of frame selection from video sequences based on bounding box area and temporal spacing. The video is divided into three segments (visualized as red lines), and one frame with the highest relative area is selected from each segment.
@@ 2.4 Dataset Annotation
All extracted video frames and images from camera traps were further processed, i.e., we manually annotated the data for three computer vision tasks: instance segmentation, individual identification, and animal pose estimation. We used Computer Vision Annotation Tool (CVAT), an open-source platform for manual labeling, whose web-based interface allows us to efficiently annotate polygons, masks, key points, and attribute tagging. The data for instance segmentation were annotated using a semi-automated approach based on the Segment Anything Model (SAM) [30]. Human annotators used positive and negative points to prompt SAM to generate initial masks, which were (if needed) manually fixed in CVAT to outline each lynx accurately.
The data targeted for animal pose estimation were also annotated using a semi-automated approach. Initial keypoint predictions were generated with a pre-trained AnimalPose model [31] and subsequently updated manually in CVAT if needed. Up to 20 key points were assigned to each individual based on the MMPose animal skeleton specification, covering joints, facial features, and the tail base. The detailed description of all keypoints and sample poses is described in Table 3 and Figure 2. In cases of partial visibility, some key points were left unannotated if they were not visible in the image. In addition to spatial labels, we also annotate image descriptive metadata, including the lynx’s viewpoint (e.g., left flank, right flank, frontal, or rear), an identifiability flag indicating whether individual recognition is possible.
Keypoint No. 1 2 3 4 5 6 7 8 9 10
[…]
Keypoint No. 11 12 13 14 15 16 17 18 19 20
[…]
Table 3: Keypoint definitions used for pose estimation.
Figure 2: Lynx lynx pose annotations. 20-keypoint skeletal structure based on AnimalPose standard.

### v2

@@ Abstract
We introduce CzechLynx, the first large-scale, open-access dataset for individual identification, pose estimation, and instance segmentation of the Eurasian lynx (Lynx lynx). CzechLynx contains 39,760 camera trap images annotated with segmentation masks, identity labels, and 20-point skeletons and covers 319 unique individuals across 15 years of systematic monitoring in two geographically distinct regions: southwest Bohemia and the Western Carpathians. In addition to the real camera trap data, we provide a large complementary set of photorealistic synthetic images and a Unity-based generation pipeline with diffusion-based text-to-texture modeling, capable of producing arbitrarily large amounts of synthetic data spanning diverse environments, poses, and coat-pattern variations. To enable systematic testing across realistic ecological scenarios, we define three complementary evaluation protocols: (i) geo-aware, (ii) time-aware open-set, and (iii) time-aware closed-set, covering cross-regional and long-term monitoring settings. With the provided resources, CzechLynx offers a unique, flexible benchmark for robust evaluation of computer vision and machine learning models across realistic ecological scenarios.
@@ 1 Background & Summary
[…]
Straka et al. [12] used a subset of the CzechLynx dataset as a case study for 2D pose estimation of endangered species. They treated Eurasian lynx camera trap images as a testbed to compare pre-trained backbones, fine-tuning strategies, data augmentation schemes, and different mixes of real and synthetic training data, using HRNet-w32 [13] and related architectures as base models. Their experiments reveiled how real and synthetic data can be combined to support pose estimation under data-limited conditions.
@@
Picek et al. [14] used the part of the CzechLynx dataset from Friends of the Earth Czech Republic (FoE CZ) to evaluate a method for individual animal identification, which models foreground and background information separately. They focused on the challenges of identifying lynx individuals over time and across different locations, using data from the 15 years of camera trap monitoring in Central Europe. The dataset enabled them to demonstrate how incorporating spatial and temporal priors, along with their Per-Instance Temperature Scaling (PITS), improves identification accuracy, particularly when animals appear in new areas or exhibit changes in appearance over time.
@@
Dula et al. [8] and Palmero et al. [9] used some parts of the CzechLynx dataset to estimate population densities by spatial capture-recapture models and other demographic parameters of the Eurasian lynx in the West Carpathians in the Czech-Slovakia borderland and in the Bohemian Forest Ecosystem: Bavarian Forest National Park in Germany and the Šumava National Park in the Czech Republic. By identifying lynx individuals, Dula et al. [8] estimated apparent survival, transition rates, and population turnover. Fluctuating densities and high turnover rates indicated human-caused mortality, which could limit population growth and the dispersal of lynx to other adjacent areas, thus undermining the favourable conservation status of the Carpathian population.
Adam et al. [15, 16] used the CzechLynx dataset as part of the AnimalCLEF 2025 competition on individual animal identification. Their goal was to establish a standardized benchmark for evaluating algorithms under realistic open-set conditions across multiple species. The CzechLynx dataset was chosen because its scale, temporal depth, and natural visual diversity provide a realistic and challenging testbed for evaluating the generalization and robustness of identification models. Its use in the challenge demonstrated the dataset’s suitability for large-scale benchmarking and its potential to support community-driven development of transferable wildlife monitoring methods.
@@ 1.1 Related Work
Animal-focused image datasets are now well cataloged and serve as a central resource for developing and evaluating methods for individual identification of animals across species [17]. Available resources cover a wide range of taxa, including large terrestrial mammals (e.g., zebras, elephants), marine species (e.g., belugas, whale sharks, turtles), and carnivores (e.g., leopards, hyenas, tigers). However, geographical coverage is not evenly distributed: most publicly available datasets originate from North America and Africa, while European resources remain relatively scarce, in part because of data privacy and conservation constraints. Large carnivores on EU red lists are particularly underrepresented in open, image-based identification datasets. A summary of non-public datasets is provided in Table 1, and public datasets with their key statistics are listed in Table 2. Short descriptions of selected public datasets are provided below.
@@
Table 1: Non-public animal re-id datasets with large carnivores.
The ATRW dataset [26] contains 8,076 videos collected from ten zoos in China and covers 92 Amur tigers (Panthera tigris). Each image is annotated with bounding boxes, identity labels, and pose keypoints, allowing development of re-identification and pose estimation methods for controlled environments and populations.
The Hyena ID 2022 and Leopard ID 2022 datasets [27], developed by the Botswana Predator Conservation Trust, contain 3,129 images of 256 individual spotted hyenas (Crocuta crocuta) and 6,795 images of 430 African leopards (Panthera pardus), respectively. Images were taken in natural settings and annotated with bounding boxes and viewpoints, making these datasets valuable for testing in-the-wild generalization performance for identification of individual animals.
The LionData dataset [28] was compiled as part of the Mara Masai project in Kenya and includes 750 images of 98 lions (Panthera leo). Each image was manually annotated to support training and evaluation of animal re-identification models under natural variability, such as lighting changes and different camera angles.
@@
Name Species Records Individuals Span BBoxes Masks Pose Wild
@@
ATRW [26] Pantheria tigris 8,076  92 1 day $\bullet$ $\bullet$ $\bullet$ $\circ$
(our) CzechLynx [30] Lynx lynx 39,760 319 15 years $\bullet$ $\bullet$ $\bullet$ $\bullet$
Table 2: All publicly available datasets with large carnivores. For each dataset, we report the number of images, unique individual identities (Ids), temporal span (Span), and the availability of bounding boxes (BBoxes), segmentation masks (Masks), pose keypoints (Pose), and in-the-wild imagery (Wild). A filled bullet ($\bullet$) indicates that the annotation type is provided, while an open circle ($\circ$) indicates it is not. Datasets derived from video recordings. Statistics for CzechLynx exclude synthetic images.
@@ 2 Methods
This section describes how the CzechLynx dataset was built, including detailed descriptions of data acquisition, annotation, and pre-processing steps. The dataset combines data collected from long-term camera trapping surveys across two distinct regions with synthetic samples generated to enhance variability. Where applicable, previously published methods are cited, and only novel technical processes are described in detail.
@@ 2.1 Dataset Collection
The CzechLynx dataset [30] comprises data collected from two independent conservation projects operating within distinct regions of the Eurasian lynx’s (Lynx lynx) Central European range. Each project used its own field methodology and data management pipeline. To maintain transparency and reproducibility, all data sources are described separately below.
@@
Carnivore Conservation Programme, Friends of the Earth Czech Republic (FoE CZ), and Department of Forest Ecology, Mendel University in Brno: This source includes data collected through long-term monitoring in the Western Carpathians (since 2009) and Southwest Bohemia (since 2015); labeled as “FoE CZ – The Western Carpathians” and “FoE CZ – Southwest Bohemia” in the following text. Camera traps were deployed on forest roads, trails, mountain ridges, as well as marking and resting sites, identified via snow-tracking, signs of occurrence, or telemetry research [31, 2]. From 2015 onward, the Western Carpathians area followed a systematic camera trapping and spatial capture-recapture model to estimate lynx densities and evaluate density fluctuations, apparent survival, transition rate, and individual turnover during five consecutive seasons [8]. Earlier data and those collected by opportunistic camera trapping allowed for estimation of minimum population size, social structure, and evidence of reproduction. Cameras used included both white-flash and infrared models (video-enabled). Most individuals were manually identified based on distinctive coat patterns visible on their flanks, forelimbs, and hindlimbs [8].
@@
Šumava National Park Administration: This source includes data collected from the areas of Šumava National Park and the Protected Landscape Area in southwestern Bohemia. The data were gathered by the Šumava National Park Administration within two long-term international monitoring projects. The first project, started in 2009, covers roughly the northwestern two-thirds of Šumava National Park. Its main goal is to obtain yearly density estimates of independent lynx individuals living in the core area of the Bohemian-Bavarian-Austrian lynx population. Within the scope of the first project, data have been collected using a $\times2.7$ km grid. White-flash camera traps were installed seasonally in every second grid cell at suitable locations along forest paths, roads, and trails. Additional details about this methodology can be found in studies by Weingarth et al. [32] and Palmero et al. [33]. The second project (initiated in 2017) aims to provide as complete an overview as possible of the entire Bohemian-Bavarian-Austrian lynx population throughout its full range. Data for this project were collected year-round at various sites, including forest paths, roads, trails, and lynx scent-marking locations, using both white-flash and infrared video-enabled camera traps. These cameras were set up continuously based on a $\times10$ km grid (ETRS LAEA 5210). In each grid cell predominantly within the Šumava National Park and Protected Landscape Area, between 4 to 8 camera sites were continuously monitored throughout the year. Further information about this project is available in a separate report [34].
@@ 2.2 Raw Data Pre-processing
The pre-processing of raw data started with the manual removal of empty images, which frequently resulted from false triggers caused by wind, moving vegetation, or changing light conditions. Once these non-informative images were discarded, the remaining photographs were screened by a team of experts and trained volunteers, who classified the content by species, while removing data with humans or vehicles. Only images positively identified as containing Eurasian lynxes were used for further analysis. All pre-selected images then underwent a second review by specialists with long-term experience in lynx monitoring. Using the species’ distinctive coat patterns on the hind limbs, forelimbs, and flanks, they manually identified individual lynxes. In each study area (southwest Bohemia and the Western Carpathians), at least three well-trained local experts independently processed and cross-checked identifications in line with established camera trapping standards [8, 35]. The resulting set of expert-verified images, each linked to a specific, individually recognized lynx, formed the basis for subsequent processing.
@@ 2.3 Video Pre-processing
The CzechLynx dataset [30] contains Lynx lynx encounters collected using different camera trap models. Some devices capture still image/images, whereas others record short videos (35% of observations). Because consecutive frames within a burst or video clip are often nearly identical, using all frames would introduce substantial redundancy and would make manual annotation time-consuming. Therefore, a small number (up to three) of the most informative frames is automatically selected for each detection event (i.e., observation).
The selection was based on automated animal detection using MegaDetector [36], a YOLOv5-based model widely used in ecological research for reliably detecting any animal in camera trap imagery[37, 38, 39]. Each video frame was processed, and those with a detection confidence below 0.7 were discarded. From the remaining frames, the first and last with valid detections were identified, and the interval between them was divided into three equal parts. One frame was then selected from each segment (specifically, the frame with the largest detected bounding box) ensuring that the chosen images capture the animal prominently and are distributed across the video timeline (see Figure 1). Selected frames were used as representative samples for downstream tasks, e.g., individual identification and pose estimation. The following criteria guided the frame selection process: (i) the animal’s full body should be visible; (ii) the animal should be captured from a side view, where identifying coat patterns are most apparent; (iii) the animal should be close enough to the camera to reveal fine details without any part being cropped; and (iv) the selected frames should be temporally spaced to capture variation in pose and movement.
Figure 1: Frame selection based on bounding box area and temporal spacing. The video is divided into three parts (visualized as delimited by red lines), and one frame with the highest relative area is selected from each segment.
@@ 2.4 Dataset Annotation
All extracted images were further processed, i.e., manually annotated, to support evaluation of three computer vision and machine learning tasks: (i) instance segmentation, (ii) individual identification, and (iii) animal pose estimation. The Computer Vision Annotation Tool (CVAT), an open-source, web-based labeling platform, was used to create polygon, segmentation mask, keypoint, and attribute annotations.
The data for instance segmentation were annotated using a semi-automated, human-in-the-loop approach based on the Segment Anything Model (SAM) [40]. Annotators used positive and negative points to prompt SAM to create initial masks, which were then checked and corrected (if needed) in CVAT to ensure accurate outlines of each lynx. To better evaluate the approach, a comparison of human- and SAM-only masks with the final human-verified annotations has been done. SAM alone reached a mean Intersection over Union (IoU) of 0.96, while fully manual annotation reached an IoU of 0.87. The human-in-the-loop workflow was both faster and more accurate (30 seconds vs 5 minutes per image), as SAM helped annotators avoid small boundary errors common in manual segmentation.
The data targeted for animal pose estimation were also annotated using a semi-automated approach. Initial keypoint predictions were generated with a pre-trained AnimalPose model [41] and subsequently updated manually in CVAT if needed. Up to 20 key points were assigned to each individual based on the MMPose animal skeleton specification, covering joints, facial features, and the tail base. The detailed description of all keypoints and sample poses is provided in Figure 2 and Table 3. In cases of partial visibility, some key points were left unannotated if they were not visible in the image. In addition to spatial labels, image descriptive metadata, including the lynx’s viewpoint (e.g., left flank, right flank, frontal, or rear), are annotated.
Figure 2: Lynx lynx pose. 20-keypoint skeletal structure based on the AnimalPose [41] standard.
1 2 3 4 5 6 7 8 9 10
$_{\mathrm{\,left}} eye$$_{\mathrm{\,right}}$ $_{\mathrm{\,left}} ear$$_{\mathrm{\,right}}$ nose throat tailbase withers $_{\mathrm{\,left}}^{\mathrm{\,front}} elbow$$_{\mathrm{\,right}}^{\mathrm{\,front}}$
11 12 13 14 15 16 17 18 19 20
$_{\mathrm{\,left}}^{\mathrm{\,back}} elbow$$_{\mathrm{\,right}}^{\mathrm{\,back}}$ $_{\mathrm{\,left}}^{\mathrm{\,front}} knee$$_{\mathrm{\,right}}^{\mathrm{\,front}}$ $_{\mathrm{\,left}}^{\mathrm{\,back}} knee$$_{\mathrm{\,right}}^{\mathrm{\,back}}$ $_{\mathrm{\,left}}^{\mathrm{\,front}} paw$$_{\mathrm{\,right}}^{\mathrm{\,front}}$ $_{\mathrm{\,left}}^{\mathrm{\,back}} paw$$_{\mathrm{\,right}}^{\mathrm{\,back}}$
Table 3: Numbered anatomical definitions of the 20 keypoints used to construct the Eurasian lynx skeleton (i.e., pose), following the AnimalPose [41] standard and the MMPose animal skeleton definition.